QUESTION 2 The price of nickel has increased to the point where once-defunct nickel mines in Nkomati, Russia have been reopened because of their potential profitability. Assume that nickel sells for $1.50 per pound. Suppose the Nkomati mine has total variable costs of $21 million when it operates at 100% capacity and fixed costs of $19 million per year. Assume that 100% capacity is associated with 23,000 tons of output per year. If the mines actual production is only 19,000 tons per year, what is the mines annual profit? Assume that variable costs are a linear function of the quantitydof nickel produced, and that 1 ton is equivalent to 2,000 pounds. Report your answer to the nearest dollar
Answer #1

Answer:

Given

for 23000 tons of output variable cost=$21 millions

so for 19,000 tons of output variable cost=(19000/23000)*21=$17.34 millions

So Total cost =Variable cost+ fixed cost=17.34+19=$36.34 millions

Selling price=$1.5 per pound

Revenue from sell of 19000 tons pf nickel=19000*2000*1.5=$57 millions

Annual Profit = Revenue - Total Cost=57-36.34=$20.66 millions